OpenStudy (anonymous):

If cscx=3 and tanx<0, find sinx, cosx, and tanx

OpenStudy (anonymous):

This is what I have so far: cscx=1/sinx 3=1/sinx 3=1/(1/3) sinx=1/3 I can't figure out how to get the cosx but I think once I get cosx, that tanx will be easy since it's just sinx/cosx

OpenStudy (johnweldon1993):

You can always use the fact that \(\large sin^2(x) + cos^2(x) = 1\)

Nnesha (nnesha):

use the hint given sin is positive and tan x is less than 0 sign chart (CAST sign chart)